# Parallel Collisionless Shocks in strongly Magnetized Electron-Ion Plasma. I. Temperature anisotropies

### v1

@@ 4.4 Temperature across shock transition
Figure 5: The spatial profiles for parallel (top) and perpendicular (y-component; bottom) temperatures in various simulations (color-coded) at $t\omega_{i}=9400$ (the same time used in Figure 1). In all cases, the ion temperature is represented by dashed curves, while the electron temperature is represented by solid curves with dots indicating the locations of bin centers where these temperatures are computed, and use an equally spaced bins that extend over $40\,c/\omega_{i}$ in all cases.
@@
where $m_{s}$ is the species (ions or electrons) rest mass, the direction $j=x$ or $y$ or $z$, and $\gamma$ is the lorentz factor. The averaging in Equation (14) is over all particles in some region of the simulation where the temperature is measured.
In the presence of non-thermal particles, this approach, however, could significantly overestimates the actual temperature of the thermal bulk of the distribution. Therefore, we implement an iterative process where we determine $\left\langle u_{j}\right\rangle$ and $T_{j}$, and then include only particles where, $u_{j}\in\left\langle u_{j}\right\rangle+3\times[\sqrt{T_{j}},-\sqrt{T_{j}}]$. We, then, iteratively recompute $\left\langle u_{j}\right\rangle$ and $T_{j}$ from the filtered data. This procedure converges typically after 3-5 times iterations giving excellent agreement with the temperatures found from fitting the full particle distribution [^1] We chose the interval $3\times[\sqrt{T_{j}},-\sqrt{T_{j}}]$ so that, once non-thermal particles are excluded, more than 99 % of the thermal particles are still incorporated. Increasing the pre-factor results in slower convergence..
@@
The Top and bottom panels in Figure 5 show the spatial profiles for parallel, $T_{x}(x)$, and perpendicular, $T_{y}(x)$, respectively, where various simulations are shown with different colors. This is shown for both ions and electrons separately: ion temperatures are indicated by dashed lines, while electron temperatures are shown as solid lines. Dots highlight the locations of bin centers where these temperatures are computed. We use an equally spaced bins when computing the temperatures that extend over $40\,c/\omega_{i}$ in all cases.
@@
The top (bottom) panel of Figure 5 displays the spatial profiles of the parallel (perpendicular; y-component) temperatures for selected simulations. Ion temperatures are represented by dashed lines, while electron temperatures are shown with solid lines. The top panel reveals a jump in the parallel temperature across all simulations, whereas the bottom panel indicates that such a jump only occurs in simulations with $\sigma_{i}<1$. In contrast, both electron and ion temperatures are well conserved for strongly magnetized shocks with $\sigma_{i}\geq 1$.
@@ 4.5 Pressure anisotropy
The results reveal a prominent jump in both ion and electron (parallel) plasma-$\beta$ across all simulations, regardless of magnetization level. In the downstream region of strongly magnetized shocks ($\sigma_{i}\geq 1$), the level of anisotropy is significantly higher compared to simulations with lower magnetization. However, strongly magnetized shocks exhibit a much smaller parallel plasma-$\beta$ for both electrons and ions.
As shown in Panel (c), electrons in all simulations remain within the stability regime, indicating that the anisotropy in electron temperatures does not drive mirror or firehose wave modes unstable. Specifically, the spatial profile of $\beta_{x}(A_{e}-1)$ for electrons always lies within the stability region, denoted by the gray area in Panel (c).
@@ 5 Conclusions
In this paper, we investigate the impact of strong magnetic fields on collisionless electron-ion shocks using Particle-in-Cell (PIC) simulations. We explore how varying the ion magnetization parameter $\sigma_{i}$ affects shock jump conditions, density, entropy, and temperatures.
@@
  • Density Compression: Strongly magnetized shocks ($\sigma_{i}>1$) exhibit a lower compression ratio ($R\sim 2$), while weakly magnetized shocks show $R>2$. At low magnetization, downstream density variations indicate simulations have not reached steady state.
  • Entropy increase: An entropy jump occurs in all shocks, with smaller magnitudes in strongly magnetized simulations ($\sigma_{i}\geq 1$) for both electrons and ions.
  • Particle Acceleration: In weakly magnetized shocks ($\sigma_{i}<1$), ions and electrons are heated to a supra-thermal population, with most upstream kinetic energy thermalized downstream. In strongly magnetized shocks ($\sigma_{i}=4$ and $10$), electron acceleration is absent, and full electron thermalization occurs downstream. Ion thermalization remains incomplete at $t\omega_{i}=9400$.
@@
  • Plasma-$\beta$ and Instabilities: A jump in ion and electron (parallel) plasma-$\beta$ is observed in all simulations. Weakly magnetized shocks ($\sigma_{i}<1$) drive mirror and/or firehose ion-unstable wave modes both upstream and downstream, while in strongly magnetized cases ($\sigma_{i}\geq 1$) suppress these instabilities in both species.

### v2

@@ 4.4 Temperature across shock transition
Figure 5: The spatial profiles for parallel (top) and perpendicular (y-component; bottom) temperatures in various simulations (color-coded) at $t\omega_{i}=9400$ (the same time used in Figure 1). In all cases, the ion temperature is represented by dashed curves, while the electron temperature is represented by solid curves with dots indicating the locations of bin centers where these temperatures are computed, and we use an equally spaced bins that extend over $40\,c/\omega_{i}$.
@@
where $m_{s}$ is the species (ions or electrons) rest mass and the direction $j=x$ or $y$ or $z$. The averaging in Equation (14) is over all particles in some region of the simulation where the temperature is measured.
In the presence of non-thermal particles, this approach, however, could significantly overestimate the actual temperature of the thermal bulk of the distribution. Therefore, we implement an iterative process where we determine $\left\langle u_{j}\right\rangle$ and $T_{j}$, and then include only particles where, $u_{j}\in\left\langle u_{j}\right\rangle+3\times[\sqrt{T_{j}},-\sqrt{T_{j}}]$. We then iteratively recompute $\left\langle u_{j}\right\rangle$ and $T_{j}$ from the filtered data. This procedure typically converges after 3-5 times iterations giving excellent agreement with the temperatures found from fitting the full particle distribution [^3] We chose the interval $3\times[\sqrt{T_{j}},-\sqrt{T_{j}}]$ so that, once non-thermal particles are excluded, more than 99 % of the thermal particles are still incorporated. Increasing the pre-factor results in slower convergence..
@@
The top and bottom panels in Figure 5 show the spatial profiles for parallel, $T_{x}(x)$, and perpendicular, $T_{y}(x)$, respectively, where various simulations are shown with different colors for both ions and electrons separately: ion temperatures are indicated by dashed lines, while electron temperatures are shown as solid lines. Dots highlight the locations of bin centers where these temperatures are computed. We use equally spaced bins when computing the temperatures that extend over $40\,c/\omega_{i}$ in all cases.
@@
The top panel of Figure 5 reveals a jump in the parallel temperature across all simulations, whereas the bottom panel indicates that such a jump only occurs in simulations with $\sigma_{i}<1$. In contrast, both electron and ion temperatures are well conserved for strongly magnetized shocks with $\sigma_{i}\geq 1$ (the black and pink lines overlap).
@@ 4.5 Temperature anisotropy
The results reveal a prominent jump in both ion and electron (parallel) plasma-$\beta$ across all simulations, regardless of magnetization level. In the downstream region of strongly magnetized shocks ($\sigma_{i}\geq 1$), the level of anisotropy is significantly higher, i.e., $A=T_{\perp}/T_{\parallel}\ll 1$, compared to simulations with lower magnetization. However, strongly magnetized cases exhibit a much smaller parallel plasma-$\beta$ for both electrons and ions.
When considering instabilities driven by ion and electron temperature anisotropies in our simulations, we are restricted to instabilities in parallel electromagnetic wave modes only, as analyzed in, e.g., Sections 7.2 and 7.3 of (Gary, 1993). We calculate spatial profiles of $\beta_{x}(A-1)$ for both ions and electrons (Figure 6, panels c and d). Regions satisfying $\beta_{x}(A-1)>1$ are mirror unstable444 While temperature anisotropies with $A>1$ can excite oblique mirror modes (typically the fastest-growing electromagnetic instabilities are oblique), parallel modes remain present as subdominant features (Gary, 1993). In our simulations, we are necessarily restricted to these subdominant parallel mirror modes., while that satisfy $\beta_{x}(A-1)<-1$ are firehose unstable (Gary, 1993; Bale et al., 2009). Regions where $1\leq\beta_{x}(A-1)\leq 1$ are likely stable and are shown as the gray region in panels (c) and (d) of Figure 6.
As shown in Panel (c), electrons in all simulations remain within the stability regime, indicating that the anisotropy in electron temperatures does not drive mirror or firehose wave modes unstable. The spatial profile of $\beta_{x}(A_{e}-1)$ for electrons always lies within the stability region, denoted by the gray area in Panel (c).
@@ 5 Conclusions
In this paper, we investigate the impact of strong magnetic fields on collisionless electron-ion shocks using Particle-in-Cell (PIC) simulations. We explore how varying the ion magnetization parameter $\sigma_{i}$ affects shock jump conditions, for density, entropy, and temperatures separately for both species.
@@
  • Density Compression: Strongly magnetized cases ($\sigma_{i}>1$) exhibit a lower compression ratio ($R\sim 2$), while weakly magnetized shocks show $R>2$. At low magnetization, downstream density variations (large-amplitude fluctuations) indicate simulations have not reached steady state.
  • Entropy increase: An entropy jump occurs in all cases, with smaller magnitudes in strongly magnetized simulations ($\sigma_{i}\geq 1$) for both electrons and ions.
  • Particle Acceleration: In weakly magnetized shocks ($\sigma_{i}<1$), ions and electrons are heated to a supra-thermal population, with most upstream kinetic energy thermalized downstream. In strongly magnetized cases ($\sigma_{i}=4$ and $10$), electron acceleration is absent, and full electron thermalization occurs downstream. Ion thermalization remains incomplete at $t\omega_{i}=9400$.
@@
  • Plasma-$\beta$ and Instabilities: A jump in ion and electron (parallel) plasma-$\beta$ is observed in all simulations. Weakly magnetized shocks ($\sigma_{i}<1$) drive mirror and/or firehose ion-unstable wave modes both upstream and downstream, whereas strongly magnetized cases ($\sigma_{i}\geq 1$) suppress these instabilities in both species.